Fermi National Laboratory
 

Accelerator Update for September 28, 2007 - October 5, 2007

 

Fermilab Shutdown Report
9/28/07 – 10/05/07

LINAC

Things Done This Week

  • From Recovered Power Outage
  • Ran 116 MeV Test beam To Verify LRF5 Operation
  • Brought All Systems Back Online
  • Ran 400 MeV Beam to Dump In Preparation For Booster Startup

Things to be done next week

  • Optimize Beam For Booster Operation

Booster

Work performed: Week 9 (Sep 28 th – Oct 5 th):

  • Start up activities thru all the week
    • penetrations sealed
    • walkthroughs and clean up
    • tunnel re-cored to operational keys
    • safety system and CDC tests performed
  • Power supplies commissioning started after the power outage 10/3
    • problems were addressed as they were appearing
    • continued with the new corrector power supply burn-in
      • playing commissioning tables at 15 Hz
    • RF system is up and running
  • Pursuing beam sign off today in anticipation of extracting beam to the MI8 dump over the weekend once the MI LCW resistivity has improved
    • entertaining the idea of accelerating low intensity beam without extracting in order to check system integrity

Antiproton Source

Mechanical

  • Leak Check (4)
    • Finished D/A, A40, A30, AP3
      • D:ESEP, A30 IP fixed; AP3 Pump
    • D:ISEP Shielding
    • Ion Chamber removal
    • Deb DCCT motorized stand
    • D:EKIK ceramic tube
      • Aligned; Vacuum tight after leak fix; Electrically connected; Tested
    • AP2 Water Hose Replacement
    • Leaking LQA replaced
    • Replaced leaking Deb Sextuple
    • SMA water leak; replaced
    • 4 water leaks fixed
    • A10 Obstruction Search
      • Components shuffled; Aligned; Baking
    • Brace Skew Quad Coils
    • Install new skew quad
      • Placed

Target Station

  • Water Systems (2)
  • Temperature Monitoring (2)
  • Target Rotation
  • Storage
  • SEM Module
  • LVDTs
  • Sample Retrived
  • Lens motion repair
    • Not successful
  • Air Exhaust Stack
  • 10mm-2 into transformer
  • 10mm-3 into transformer
    • On Test Stand
  • Lens replacement
    • 10mm-2 now in vault and being tested
  • HiPot of PMAG
  • PMAG water system flush
  • Lens Module water system work
  • Power/Emergency Generator Switch
  • Elevator table repair

Cryo

  • Valve Stem replace
  • Leak Check (1 of 2)
  • LN2 transfer line maintenance

Cooling Systems

  • Deb 3&4 Trans Notch Filters
  • Deb Double Notch Filters
  • Core 4-8GHz Equalizer
  • TWT balance & trip points
  • Stacktail Filter 3 cable change

Electrical (Pbar & EE Support)

  • Shunt Move
  • Shunt Repairs
  • PS Maintenance
  • Lens SCRs
  • Regulation Electronics
    • Done except A:LQ
  • Ion Pump maintenance
  • Deb Shunt re-arrangement
  • Deb Extraction kicker load adjustment

Other

  • Interlocks/Safety System
    • Vault key tree done; testing to do
  • Repair cathodic protection
  • Holes under buildings
    • Water dripping from tunnel ceiling
      • Contractor sealed ceiling
    • Await contract for filling holes
  • DCCT calibration
  • Baking System checkout
  • RF & TWT filters maintenance
  • Lights checkout & repaired
  • Acc BPM repairs
  • Activation Survey
  • Inductance checkout
  • PM ODH Fans & Sumps
  • Tunnel Wireless
  • Install AP2 Leak Detector
  • SEM maintenance
    • In progress
  • Clogged sump line – Transport
    • Berm dug; Replaced entire line; re-established berm; to be surveyed

To Re-establish Operations

  • Reverse Protons
    • {1 shift} re-establish MI-ACC ($2D/30sec)
    • {3+ shifts} Accumulator (1-shot timeline)
      • BPM checkout; Orbit Checkout; Aperture (A10, running wave); Tunes across aperture; Calibrate emittance monitors
    • {1+ shifts} Debuncher (1-shot timeline)
      • Establish beam; Orbit Checkout; Aperture (EKIK, running wave)
  • Stacking
    • {1+ shifts} 120GeV operations ($29/4.4sec)
      • Protons to Target; Secondaries to Deb (include 120GeV orbit for lens height); Kickers & RF systems timing; pbars to core
    • {2 shifts} Check phasing of cooling systems
    • 20mA {<0.5 shift} phase dampers
    • 40mA to 60mA transfer to Recycler

Further Studies

  • Clean Accumulator Studies [Reverse Protons]
    • {1 shift} D/A optics ($2D/min)
      • Deb Rev Prot TBT; D/A BPMs?; SEMs
    • {1+ shift} D/A Transfer Efficiency ($2D/min)
      • Set Acc transfer scrapers to define different beam sizes
        • Record DCCTs, SEMs, Loss monitors
    • {1+ shift} A10 & running wave (1 shot timeline)
      • Make component limiting aperture and perform running wave
        • Do different limiting apertures cause running wave response?
    • {1 shift} Accumulator optics – all orbits (1 shot timeline)
    • {1 shift} Commission SQ107 (1 shot timeline)
      • Includes coupling correction and measuring vertical dispersion
  • Accumulator hold onto core [Reverse Protons]
    • {1 shifts} Deb Optics (1 shot timeline)
      • Measure; Change; Measure
  • Stacking
    • {0.5 shift} Commission Debuncher switch-in filters
    • {1+ shift} D/A Transfer Efficiency (long cycle time; consistent beam)
      • Vary Debuncher cooling time
        • Record IPM, Loss Monitors, production rate

Main Injector

Activities this week

  • MI52 Beam tube replacement
    • Beam pipe replaced
    • Lambertson bake underway
  • MI40 beam tube replacement
    • Bake out complete
  • MI30 collimators
    • LCW Hose installed
    • Collimator survey / alignment finished
    • Concrete wall complete
  • LCW
    • Should be sufficiently resistive by mid-week

MI Startup Activities

  • General Startup Plan
    • Turn on to pre-shutdown conditions.
      • Slip stacked beam to Pbar target
      • 2+5 mode for Numi and Mixed mode
      • Use the old damper system
      • Use the pre-shutdown TLG modules
    • 2.2s cycle spacing
    • We would like to switch to multibatch slip stacking in 2 weeks
      • Need new TLG modules and commissioning activities must be complete.
  • Measurements requiring dedicated time
    • Establish 8 Gev. beam. 4 hrs.
    • MI8 line lattice measurements and matching 2 x 4 hrs.
    • Establish 120 Gev. beam. Smooth orbits . 4 hrs.
    • Aperture scans at MI30 2 hrs.
    • Quad moves??? 8 hrs.
    • Coalescing tune up 4 hrs.
    • Slip Stacking tune up. 2 hrs..
    • Tune up extraction positions and abort line 4 hrs. per line
  • “Parasitic” Measurements
    • Commission new damper.
    • Collimator commissioning (MI8 and MI30).
    • Ring aperture scans with new software program.
    • General tune up.

Recycler

Recyler Work List

Not yet finished…

  • QT609 wiring (need lead blankets)
  • BLM for scrapers: rebuilding in progress (50% done)
  • BPM calibration check (40% done)
  • Elevation survey after collimator installation
  • Field check for new quadrupoles
  • LEP dipole field centering on beampipes
  • Ring wide corrector check
  • Kicker timing verification
  • RF power and correction system check (next week)
  • Safety sign-off in progress

Finished

  • New DCCT installed, baking finished
    • Observed 10E10 difference for 75 degrees C increase in temperature during the bake
  • V303 and H308 moved for MI collimator work (done)
  • Old DCCT support (done)
  • STC tank 7 leak check finished, nothing new
  • STC photodiodes alignment (done)
  • Sublimation of TSP’s from 204 to 210 (done)
  • Bulk supply upgrade (done)
  • The 800 line leak repair (done)
    • MW flange seal (fixed)
    • HP803 leak (fixed)
  • Diagnostics maintenance
    • New FW paddles (done)
    • Schottky (1.75 GHz) system gain check (done)
    • Schottky (21 MHz) L.O. check (done)
  • Trombone installation (done)
  • QCL cable positioning from 304 to 309 (done)
  • Magnet moves at 305 and 307 (done)
  • TOR330 repair (done)
  • Trombone wiring 601-608 (done)
  • Ecool study plan ready
  • End-of-shutdown checklist ready
  • Recycler Start-up plan ready

Tevatron

Vacuum Issues

  • C0 still leak checking
  • Bad dipole at A14-2 being replaced This is the 2nd bad dipole so far
  • Out side air leak at E1is fixed
  • Ready to do conning towers at E-1
  • Air or N2 leak at A2 still under investigation

SwitchYard

Continuing with maintenance and vacuum work

NuMI

NuMI Shutdown Activities

  • MI65 emergency power transfer switch repair
    • New controller on order through FESS-OPS
  • Target, Horn 1 & Horn 2 maintenance and upgrades
    • Re-installed and surveyed work cell door
    • Surveyed the target and Horn1
    • Completed as-found survey of Horn 2
    • Installed RAW water lines for Horn 1
    • Weld all weldolets on Horn1 supply lines for new pressure gauges
  • Horn 2 suction line repair
    • Completed pre-ALARA meeting with Rad Safety
    • Installed temporary shielding wall for Horn 2 repair
    • Move Horn 2 into work cell for repair on Sunday
    • Practice Horn 2 repair procedure on Saturday
    • Complete ALARA meeting with repair team on Saturday morning
  • Decay pipe window and helium line installation
    • All 1” SS helium fill pipe installed in the MINOS service building, absorber passageway & absorber hall
    • Held two coordination meetings in the past week
    • Continue to develop ODH analysis
    • Continue to develop DP helium fill system engineering notes
    • Continue to NuMI shielding assessment addendum
    • Continue to develop NuMI note and SAD
  • Target pile dehumidification installation project (80% Complete)
    • Installed dehumidifiers on to platform on east side of work cell
    • Started installing 26” & 16” round duct on dehumidifiers
    • Installed strainer for new chiller system
    • Installed condensate tank & secondary containment tank
    • Installing 4”chilled water piping in dehumidifier area
  • Target pile dehumidification installation project
    • Install electrical for dehumidifier controls in support room
    • Install electrical raceway for chillers in SB & Mezz. area
    • Installed epoxy coating on evaporator containment area
  • Upgraded of muon monitor gas system
    • Installed new bubblers out side of alcoves to reduce back diffusion of exhaust gas
    • Install calibration chambers
  • Install Experiment T968 in muon alcove 2
    • T2K muon monitor prototype
  • Experiment T972 installed in muon alcoves
    • Foils installed in muon alcoves
  • Seal MINOS Hall roof for MINERvA experiment (20% Complete)

MiniBooNE

  • Change RAW DI Bottle (by next Tuesday)
  • Run power supplies in evening (when LCW is polished).
  • Test horn. (week of Oct. 8)
  • Rad Safety walk-thru MI12 A (today)

MuCool Beamline Installation

  • Mechanical installation
    • Beam stop #1 installed
      • Assembly of beam stop #2 in progress
  • Shielding installation
    • Shielding around the beam pipe thru the concrete wall was complete early in the week
      • Last shielding requirement for beam operation in Linac
    • Linac and MTA enclosure interlocks are back to normal configuration
  • Alignment
    • Another rough alignment was performed early in the week (Linac side)
      • Necessary step in order to continue progress with vacuum connections
    • Work continue with the referencing of MW system.
  • Vacuum work
    • MTA-Linac side
      • Majority of vacuum connections were complete by Monday
        • Outstanding job done by MSD and PPD techs
      • Initial rough vacuum on Tuesday! Milestone accomplished
        • Only 2 connections had to be repaired
        • More time is needed in order to validate the vacuum in the line.
      • 2 out of 4 MW system still need to be fully installed
    • MTA-MuCool side
      • Installation of the available spool pieces resumed 10/03
        • ½ dozen adaptor flanges had to be made
      • Final spool pieces are being welded and leak checked
      • MW5-6-7 cans were delivered
        • 2 were mounted on their stands by the end of the day
  • Electrical Work
    • Instrumentation group connected all BPM and BLM cables (Linac side)
    • Awaiting for the completion of the vacuum installation to continue at MuCool side
  • Next week
    • If MuCool RF program schedule permits, the beam line installation at the MuCool side continues throughout the week
  • Plans for the future
    • A list of remaining tasks has been compiled
      • it will be submitted to the worklist database in search for an opportunity to access Linac

Collider Detector at Fermilab

  • silicon leak repair
    • good flow established on both sides
    • silicon checkout will be completed this morning
  • detector survey continues on east side
  • survey/sili checkout ®bolting down the plug ®solenoid powering
  • flammable gas flow to all detectors starts Monday ®wire chamber checkout (need solenoid powered for tracking chamber)
  • much (and most) other work completed

D-Zero

Work continues

Alignment

  • Tevatron Task Manager: Bocean Field Crew: Bradford,Teafoe, Volk.
  • 5496 Re align low beta quads and separators at B0.Date:10/05/2007 1 day 8 man crew
  • Main Injector Task Manager: Oshinowo Field Crew: Wilson, Bradford.
  • No Request
  • P-Bar Task Manager: Oshinowo Field Crew: Wilson, O’Boyle, Wyatt.
  • No Request
  • Numi Task Manager: Bocean Field Crew: Crutcher,Bradford,Wish.
  • 5584 Target, Horn 1, Horn 2, as found also old target check. Date: 10/11/2007 3 man crew
  • Reference Task Manager: Oshinowo Field Crew: Wilson, Contract Crew.
  • Reference multiwire MTA
  • Establish alignment Fiducial on 2 new MTA C-magnets, LCA001 (short) and LCB001 (long).
  • MTA Task Manager: Oshinowo Field Crew: Wilson, Contract Crew.
  • No Request
  • CDF Task Manager: Oshinowo Field Crew: Wilson, Contract Crew.
  • No Request
  • Recycler Task Manager: Oshinowo Field Crew: Wilson, Contract Crew.
  • No Request
  • 400Mev Task Manager: Oshinowo Field Crew: Wilson, Contract Crew.
  • No Request
  • Booster Task Manager:Oshinowo Field Crew:
  • No Request
  • D0 Task Manager: Oshinowo Field Crew: O’Boyle,Adkins,Wish / Wilson, J Wilson
  • "5523 Final Stickmic closing survey of muon system including CF, EF after detector is closed. This date is preliminary. To be confirmed in late September"
  • 5586 V-STAR measurements of positions of south B-layer pixel counters before B-layer walls are opened and after they have been closed. One week or less
  • Nova Task Manager: Oshinowo Field Crew: O’Boyle,Adkins,Wish / Wilson, J Wilson
  • 5477-Stake out new Recycler RR30 quad positions for the Nova Project
  • 5476-Stake out new Recycler extraction line (MI30) for the Nova Project Accelerator upgrade.
  • Civil Task Manager: Coppola Field Crew:
  • Resurvey of AP 2/3 Enclosure Berm just south of Indian Rd. recently disturbed and backfilled for repair of broken sump discharge line. Date: 9/29/2007 COMPLETED more dirt needed
  • 5611 Switchyard Enclosure C north outside stairwell surface dirt is being graded for improved drainage. Please survey 100' north and south of drainage area rework for needed shielding
  • CAF

5580 CM1 Cavity String Assembly Phase-II: rough alignment at MP9.

5581 CM1 Cavity String Assembly Phase-II: precise alignment at ICB.

5582 CM1 Cavity String testing at ICB of stability and deformation after transportation.

5583 CM1 Cold Mass and Final Assembly alignment and referencing at ICB.

  • Recycler Task Manager: Oshinowo Field Crew:
  • No Request
  • Linac Task Manager: Oshinowo Field Crew:
  • No Request
  • P150 Task Manager: Oshinowo Field Crew:
  • No Request
  • ILC Task Manager: Bocean
  • 5579 Alignment support For Beam line installation

Alignment and Metrology Home page

  • http://ppd.fnal.gov/align/align.html

Mechanical

Work in the tunnels continues.

Cryogenic System

Current Operations

  • Tevatron at LN2 temperature except A1 and E1 at room temperature.
    • Tunnel magnet cryostat vacuum at A2 being investigated by AD/MSD.
    • FRIG system has been purified using Mobile Purifier at BA, thus all Mycom comprs have been operated during purification process, as well as wet expanders and cold compressors per SOP.
    • FRIG D2 helium-atm leak on low pressure heat exchanger end-plate was found and cracked weld readily repaired; heat exchangers ringwide were inspected.
  • CHL helium inventory management system operating on clean Helium Tank Farm, with A0 GHe Recovery back online to CHL Mobile Purifier.
  • CHL Helium Coldbox-1 purification in progress; Comprs C and D have been purified and anticipate CB-1 cooldown later today thus being cold by next week in preparation of Tevatron cooldown and liquid helium delivery from vendor.
  • CHL Nitrogen Reliquefier (NRL) plant off and LN2 supply dependent on vendor deliveries to CHL LN2 dewar tanks; deliveries are fine. NRL IR Booster and TSI comprs have been started and plant purified; anticipate cooldown of plant next week.
  • CHL 24/7 operations shiftwork resumed on Monday Oct. 1.

FRIG Maintenance in progress

  • 13 of 13 wet expander flywheels have been replaced.
  • 26 of 26 wet/dry expander overhauls of cold ends are completed (as well support for CDF & DZero expanders).
  • Various vacuum, instrumentation, and controls maintenance/repairs continuing.
  • Mycom compr startup checkout completed with 31/31 available.

CHL Maintenance in progress

  • Coldbox-1:
    • Startup activities in progress
  • NRL plant:
    • Subsystems startups in progress.
    • Planning for Niagara cooling tower cleaning if present pleasant weather conditions continue.
  • Various compr and facility maintenance in progress

Controls

  • Recovered from Wednesday Power Outage
  • Round 2 of SDA and S&R Database Updates – Complete
  • MI-30 Collimator Controls Installation & Test - Complete
  • General Maintenance

FESS

CUB

  • DDC work – control wiring/upgrades under going system checks.
  • PBar S/T heat exchanger [HE-8] is under evaluation regarding end of life. Eddy current testing shows up to 30% of its tubes are over/at/near the 50% thickness factor [15 leaking tubes were plugged]. This is the back-up exchanger for the P/F HE-9. We will be able to continue to use but plans should be made for its replacement.
  • The 55LCW pump replacements are complete [LP4 & 5]. LP5 is on line but we may require some short shutdowns to finalize automatic control configuration.

Casey’s Pond

ICW FLOW METER WIP. Transfer back to the 16 inch strainer on Mon 10/01/07 was successful.

Poles

Scheduled site outage/MSS energizing on 10/3/07 was successful.

TeV pond silt and Extension Collars

  • Completed.
  • Some concern on the condition of the TEV T&S heat exchangers due to silt/plugage which the AD would check on a routine basis.

Craft SD Status

Shift Work has Resumed starting Monday 10/01/07 with rotating shifts kicking in this weekend.

  • Electrical craft backlog remains high & has been prioritized to finish critical work.
  • Most of the Mechanical work is completed.
  • HVAC is re-checking tunnel dehumidifiers/filters.
  • FINDINGS – can review at the last meeting.

Other Jobs

  • TENTATIVE SITE ICW FLUSH started Mon 10/01/07 THRU ~10/08/07.
  • ZEBRA MUSSEL TREATMENT SCHEDULED FOR MID-OCTOBER.
  • WILSON HALL FIRE PUMP RUN [ICW] test to be scheduled by end of October to satisfy DOE requirements.
  • WILSON HALL power shutdown TBS in November.

Safety

  • Safety System Status
    • Upgrade M02 - M05 Enclosure Interlocks 100%
    • Upgrade M01 Enclosure Interlocks 100%
    • Upgrade Main Injector PLC Code and replace PLC Batteries 100%
    • Replace Pbar Vault Key Tree Door Latches & add status LEDs 100%
    • Upgrade Meson Area Oxygen Monitors 10%
    • Replace Safety System DEC talk with ATT Natural Voices 100%
    • Modify Linac Enclosure Interlocks 100%
    • Linux kernal 2.4 to 2.6 upgrade to Safety Data Acquisition System 100%
    • Route Key Tree release signals from MCR to Meson 100%
    • Install Radiation Monitor Cables for the Muon Test Area 100%
    • Replace oxygen monitor cells in the Tevatron 100%
  • Radiation Exposure
    • ~0.94 person-rem total dose for the eighth week of the shutdown.
    • ~8.4 person-rem accumulated dose for the shutdown.
  • Startup
    • NTF Safety System testing completed
    • Linac Safety System testing completed
    • Booster Safety System testing completed (except for Booster beam to Main Injector portion)
    • 8 GeV Line Safety System testing completed
    • Linac and NTF re-start sign-offs completed and Beam Permits and Running Conditions issued, allowing operation
  • A successful shutdown is getting our work done safely

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 10/05/2007    email Fermilab